OpenStudy (anonymous):

Which Property of Equality would be best to use first to solve the equation -9n + 5 = -67 for n? Addition Property of Equality Subtraction Property of Equality Multiplication Property of Equality Division Property of Equality

OpenStudy (paki):

-9n + 5 = -67 for "n" -9n+5-5=-67+5.... cancel (+5-5)... so which property i used ....... @cheeta14

OpenStudy (paki):

@cheeta14 you got the point..... which i did...